OpenStudy (anonymous):

Break-Even Analysis: A Look at Profit and Loss: As an entrepreneur, there are going to be many decisions that you need to make, such as the price to charge your customers for your goods and services. You have just graduated from college and recently opened a specialty pizza restaurant. Based on surveys conducted in your area, you determine that it is feasible to sell your specialty pizzas for $15. The cost for making the pizzas includes a fixed cost of $55 and a labor cost of $4 per pizza.

OpenStudy (anonymous):

so would i take the 15-4= 11

OpenStudy (anonymous):

to do an equation for a revenue, would i take 15-4=11

OpenStudy (anonymous):

to do an equation for a revenue,pizza price = $15 cost = ($4 * number of pizzas) + $55 revenue = (pizza price) * (number of pizzas) profit = revenue - cost break even occurs when revenue = cost